# U-MARVEL: Unveiling Key Factors for Universal Multimodal Retrieval via Embedding Learning with MLLMs

### v1

@@ 3.3 Will Reranker Distillation Improve Performance?
Some studies [36, 30] further refine retrieval results by applying a reranking step on top of the recall model. The reranker can be either a trained model [36] or a zero-shot reranker [30] based on MLLMs. While this recall-then-rerank strategy often improves retrieval performance, it introduces additional inference latency and increases system complexity. In this work, we exploit to distill this cascaded pipeline into a single unified model, thereby simplifying the system while maintaining high performance.
@@
Inspired by [36], we train a generative reranker based on the decoder-only MLLM architecture. The training data is prepared by: (1) extracting embeddings for all queries and candidates via the recall model from Sec. 3.2.2, then (2) constructing pairs for each query consisting of a positive match and the top 50 most challenging negatives. The model is prompted to output "YES" for positives and "NO" for negatives, optimized with next-token prediction loss to maximize the likelihood of correct response.
@@
After training the reranker, we employ it to refine the ranking results from the recall model, further improving the retrieval performance. Specifically, the recall model from Sec. 3.2.2 first retrieves the top- candidates, then each (query, candidate) pair is processed by the reranker sequentially. The reranker is prompted to judge each candidate with a "YES" or "NO" response, where the probability of "YES" serves as the reranker’s score. To further enhance overall performance, we combine the scores from both models through linear interpolation: , where becomes the final relevance score, and is a weight hyperparameter. This completes our recall-then-rerank pipeline, producing optimized query-candidate matching scores.
@@
We distill this entire pipeline as a teacher model into a single student model, where the combined recall-then-rerank knowledge is compressed into one unified model. Specifically, the recall model is continue trained using contrastive learning with positive and top- negative samples, optimized via KL divergence:
where represents the multi-model ensemble scores and the single model’s predictions. The scores are softmax-normalized before computing the KL divergence. The inference time complexity of the model is , making full-scale distillation computationally expensive. Notably, our distillation approach achieves significant performance improvements by distilling on only of the data, demonstrating the remarkable efficiency and effectiveness of our distillation strategy.
@@
As shown in Table 6, the "Reranker" represents the model we aim to distill, and the "Distillation" method achieves an improvement of 1.5%/0.8% over the baseline. However, "Distillation" model is built upon baseline model, with its training data generated by re-ranking the top- retrieval results from baseline model by using a "Reranker". As a result, the training data correspond to hard negative samples for the baseline model. To investigate whether the performance improvement of "Distillation" model is driven by or the effect of knowledge distillation or the introduction of hard negative samples, we designed experiment "Continue-hard" for validation. Unlike the "Distillation" stage, the "Continue-hard" model continues training by treating the training data as hard negatives. The "Continue-hard" method demonstrates an improvement of 0.5%/0.1% over the baseline; however, it exhibits a decline of 1.0%/0.7% when compared to the "Distillation" approach. It validates that the performance improvement is clearly attributable to the distillation process.
@@ 3.4 U-MARVEL Framework
Motivated by the findings presented above, we introduce a unified framework, term U-MARVEL (Universal MultimodAl RetrieVal via Embedding Learning), which consists of the following three stages: (1) Progressive transition: the model is progressively fine-tuned on the increasing complexity levels of the retrieval data, which enables the model to gradually adapt to retrieval tasks. (2) Hard Negative Mining: building on Progressive transition, the model’s retrieval performance is further enhanced through hard negative mining, accompanied by the filtering of false negatives. (3) Distillation: we perform knowledge distillation on the reranker model using the sampled dataset sequentially. Detailed experimental settings and results for U-MARVEL are provided in Table 12 in Appendix 3 and Table 13 in Appendix D, respectively.
@@ 4.1 M-BEIR Evaluation
We first compare U-MARVEL with competitors on M-BEIR benchmark [51] in both local and global pool settings, details of this dataset can be found in Appendix A. For fair comparison, we also present results of the recall-then-rerank cascade pipeline, denoted as , similar to that used in LamRA [36], details seen in Appendix 3. As shown in Table 7, our primary comparisons are conducted under the local pool setting, while detailed results for the global setting are provided in the Table 14 of Appendix D.
| Task | Dataset | Metric | Single Model | +Rerank | ||||||
CLIP-L [42] |
SigLIP [58] |
[51] |
[51] |
LamRA-Ret [36] |
U-MARVEL |
LamRA [36] |
| VisualNews | R@5 | 43.3 | 30.1 | 23.4 | 42.6 | 41.6 | 47.3 | 48 | 49.4 | |
| MSCOCO | R@5 | 61.1 | 75.7 | 79.7 | 81.1 | 81.5 | 84.4 | 85.2 | 85.6 | |
| Fashion200K | R@10 | 6.6 | 36.5 | 26.1 | 18 | 28.7 | 33.6 | 32.9 | 34.2 | |
| WebQA | R@5 | 36.2 | 39.8 | 80 | 84.7 | 86 | 97.1 | 96.7 | 98.5 | |
| EDIS | R@5 | 43.3 | 27 | 50.9 | 59.4 | 62.6 | 78.8 | 75.8 | 81.4 | |
| WebQA | R@5 | 45.1 | 43.5 | 79.8 | 78.7 | 81.2 | 88.5 | 87.7 | 89.4 | |
| VisualNews | R@5 | 41.3 | 30.8 | 22.8 | 43.1 | 39.6 | 47.3 | 48.6 | 50.5 | |
| MSCOCO | R@5 | 79 | 88.2 | 89.9 | 92.3 | 90.6 | 93.5 | 92.3 | 88.4 | |
| Fashion200K | R@10 | 7.7 | 34.2 | 28.9 | 18.3 | 30.4 | 35.1 | 36.1 | 37.7 | |
| NIGHTS | R@5 | 26.1 | 28.9 | 33 | 32 | 32.1 | 34.2 | 33.5 | 34.7 | |
| OVEN | R@5 | 24.2 | 29.7 | 41 | 45.5 | 54.1 | 62.5 | 59.2 | 63.7 | |
| InfoSeek | R@5 | 20.5 | 25.1 | 22.4 | 27.9 | 52.1 | 58.3 | 64.1 | 62.9 | |
| FashionIQ | R@10 | 7 | 14.4 | 29.2 | 24.4 | 33.2 | 36.4 | 37.8 | 38.2 | |
| CIRR | R@5 | 13.2 | 22.7 | 52.2 | 44.6 | 53.1 | 60.7 | 63.3 | 63.2 | |
| OVEN | R@5 | 38.8 | 41.7 | 55.8 | 67.6 | 76.2 | 79.4 | 79.2 | 80.8 | |
| InfoSeek | R@5 | 26.4 | 27.4 | 33 | 48.9 | 63.3 | 74.7 | 78.3 | 78.9 | |
| Avg. | 32.5 | 37.2 | 46.8 | 50.6 | 56.6 | 63.2 | 63.7 | 64.8 | ||
The proposed U-MARVEL framework establishes new state-of-the-art performance across both single-model architectures and recall-then-rerank approaches. Specifically: (1) U-MARVEL achieves superior results in 15 out of 16 retrieval tasks. On average, U-MARVEL outperforms models UniIR [51], MM-Embed [30] and LamRA-Ret [36], and demonstrates 6.6%/5.8%(local/global setting) improvement over LamRA-Ret, the previous state-of-the-art. (2) also outperforms LamRA [36] across the majority of tasks and achieves an average improvement of 1.1%/0.4%(local/global setting). (3) While LamRA requiring a two-stage recall-then-rerank inference process, U-MARVEL achieves comparable performance with only a single-stage inference. This breakthrough means that our training scheme has achieved the ranking quality that traditionally requires a cascade architecture on the single encoder model, achieved a balance between efficiency and effectiveness.
@@ 4.2 Zero-shot Evaluation
We evaluate the performance of U-MARVEL with several unseen datasets in a zero-shot manner. Details regarding the datasets and evaluation metrics are provided in Appendix B. As presented in Table 8, U-MARVEL achieves state-of-the-art performance on 6 out of 10 tasks, highlighting its exceptional zero-shot capabilities compared to leading methods such as VLM2Vec [24], UniME [15], and LamRA [36]. Furthermore, we evaluate U-MARVEL on the zero-shot text-to-video retrieval task. The results on the MSVD and MSR-VTT datasets, summarized in Table 9, demonstrate that U-MARVEL outperforms VLM2Vec [24], LamRA [36], and LLaVE-7B [25]. However, it is worth noting that exhibits degraded performance, which may be attributed to fail to model temporal contexts, such as action sequence progression.
| Task | Dataset | Metric | Single Model | +Rerank | ||||||||
CLIP-L [42] |
[51] |
E5-V [22] |
MagicLens-L [60] |
VLM2Vec [24] |
UniME [15] |
LamRA-Ret [36] |
U-MARVEL |
LamRA [36] |
| ShareGPT4V | R@1 | 84 | 85.8 | 86.7 | 85.5 | 90.7 | 97.2 | 93.3 | 96.4 | 97.9 | 97.8 | |
| R@1 | 52.8 | 75 | 84 | 59.3 | 90.8 | 95.9 | 95.1 | 96.7 | 98.8 | 97.7 | ||
| Flickr | R@1 | 67.3 | 78.7 | 79.5 | 72.5 | 76.0 | 81.9 | 82.8 | 85.4 | 88.1 | 88.5 | |
| ShareGPT4V | R@1 | 81.8 | 84.1 | 84 | 60.9 | 85.8 | 93.9 | 88.1 | 97.2 | 96.5 | 98.9 | |
| R@1 | 68.7 | 78.4 | 82.4 | 24.2 | 84.7 | 95.2 | 94.3 | 93.5 | 98 | 98.2 | ||
| Flickr | R@1 | 87.2 | 94.2 | 88.2 | 84.6 | 90.6 | 93.4 | 92.7 | 93.3 | 97.6 | 95.1 | |
| MAP@5 | 4 | 12.5 | 24.8 | 29.6 | - | - | 33.2 | 36.2 | 42.8 | 46.0 | ||
| R@1 | 13.3 | 16.8 | 18.5 | 16.3 | - | - | 18.9 | 19.1 | 24.8 | 22.6 | ||
| R@1 | 23.7 | 26.8 | 54.6 | 28 | - | - | 62.8 | 70.3 | 70.9 | 77.6 | ||
| R@5 | 17.7 | 39.4 | 19.2 | 22.6 | - | - | 60.9 | 65.7 | 63.9 | 66.3 | ||
| ITM | CC-Neg | Acc. | 66.7 | 79.9 | 83.2 | 62.7 | - | - | 79.6 | 84.5 | 85.9 | 86.1 |
| Acc. | 73 | 80.3 | 84.7 | 75.9 | 79.5 | 85.0 | 85.8 | 87.9 | 93.5 | 93.4 | ||
| * indicates that the images in these datasets are sourced from COCO or FashionIQ. | ||||||||||||
@@
| InternVideo [50] | 40.0 | 65.3 | 74.1 | 43.4 | 69.9 | 79.1 |
| ViCLIP [48] | 42.4 | - | - | 49.1 | - | - |
| UMT-L [29] | 42.6 | 64.4 | 73.1 | 49.9 | 77.7 | 85.3 |
| InternVideo2s2-6B [49] | 55.9 | 78.3 | 85.1 | 59.3 | 84.4 | 89.6 |
@@
| VLM2Vec [24] | 43.5 | 69.3 | 78.9 | 49.5 | 77.5 | 85.7 |
| LamRA [36] | 44.7 | 68.6 | 78.6 | 52.4 | 79.8 | 87.0 |
| LLaVE-7B [25] | 46.8 | 71.1 | 80.0 | 52.9 | 80.1 | 87.0 |
| U-MARVEL | 47.2 | 72.0 | 80.2 | 54.6 | 80.9 | 87.7 |

### v2

@@ 3.3 Will Reranker Distillation Improve Performance?
Some studies (Liu et al., 2024b; Lin et al., 2024) further refine retrieval results by applying a reranking step on top of the recall model. The reranker can be either a trained model (Liu et al., 2024b) or a zero-shot reranker (Lin et al., 2024) based on MLLMs. While this recall-then-rerank strategy often improves retrieval performance, it introduces additional inference latency and increases system complexity. In this work, we exploit to distill this cascaded pipeline into a single unified model, thereby simplifying the system while maintaining high performance.
@@
Inspired by (Liu et al., 2024b), we train a generative reranker based on the decoder-only MLLM architecture. The training data is prepared by: (1) extracting embeddings for all queries and candidates via the recall model from Sec. 3.2.2, then (2) constructing pairs for each query consisting of a positive match and the top-50 most challenging negatives. The model is prompted to output “YES” for positives and “NO” for negatives, optimized with next-token prediction loss to maximize the likelihood of correct response. More detailed specifics can be found in Appendix C.
@@
After training the reranker, we employ it to refine the ranking results from the recall model, further improving the retrieval performance. Specifically, the recall model from Sec. 3.2.2 first retrieves the top- candidates, then each (query, candidate) pair is processed by the reranker sequentially. The reranker is prompted to judge each candidate with a “YES” or “NO” response, where the probability of “YES” serves as the reranker’s score. To further enhance overall performance, we combine the scores from both models through linear interpolation: , where becomes the final relevance score, and is a weight hyperparameter (fixed at 0.5). This completes our recall-then-rerank pipeline, producing optimized query-candidate matching scores.
@@
We distill this entire pipeline as a teacher model into a single student model, shown in Figure 2, where the combined knowledge is compressed into one unified model by using KL divergence defined as Eq. 7.
where represents the multi-model ensemble scores and the single model’s predictions. The scores are softmax-normalized before computing the KL divergence. In traditional distillation frameworks, the base model typically aligns with the reranker. Distillation samples are structured as , with similarity matrices incorporating in-batch negatives. However, recall-then-rerank inference over this similarity score matrix incurs prohibitive computational costs, making it impractical for deployment.
We propose an improved distillation method to address this limitation. Specifically, we construct samples in the form of (query, positive, top- hard negatives), and confine distillation to the scope of positive and hard negative samples. In this way, it significantly reduces computational overhead while substantially increasing the diversity of features encountered during model training. Thus, it achieves both resource efficiency and performance gains—critically, it makes distillation practically feasible. Furthermore, besides the data and base model, a key factor contributing to the effectiveness of this distillation method lies in its tight alignment with the recall-then-rerank inference pipeline, depicted in Figure 2: (1) Retrieving with a hard negative model; (2) Reranking the top- results and generating integrated scores. Our distillation method mirrors this workflow: it adopts the hard negative model as the base and distills top- scores from the integrated model.
To further validate its efficiency and effectiveness, we provide the following mathematical analysis. Let denote the number of queries in a batch, and the total number of queries. In traditional distillation, the computational complexity of inferring the similarity matrix is , and the number of features observed per training epoch is . The total time for inference and training is . In our improved method, the cost of computing the similarity matrix is reduced to , and the number of features observed during training increases to . The total time is reduced to . Substituting our actual parameters, and , the improved method reduces the total time to of the original, while the number of features seen during training increases by . In terms of actual training time, improved distillation completes in just 14 hours, compared to a theoretical cost of over 340 hours for the traditional method. This substantial efficiency gain makes MLLM distillation practically feasible, transforming it from an intractable task to a manageable one. More detailed information can be found in Appendix E and Table 14.
@@
As shown in Table 6, the “Recall-Reranker” represents the model we aim to distill, and the “Distillation” method achieves an improvement of 1.5%/0.8% over the baseline. However, distillation data are generated from the top- retrieval results from baseline model and correspond to hard negative samples for it. We designed experiment “Continue-hard” for validation. Unlike the “Distillation” stage, the “Continue-hard” model continues training by treating the training data as hard negatives. The “Continue-hard” method demonstrates an improvement of 0.5%/0.1% over the baseline; however, it exhibits a decline of 1.0%/0.7% when compared to the “Distillation” approach. It validates that the performance improvement is clearly attributable to the distillation process. Notably, contrary to the instability observed in Section 3.2.2 with the “only top- hard negatives” setting, the “Continue-hard” strategy () employed here remains stable and achieves slight gains. This stability is primarily attributed to the strong initialization: starting from a converged Recall model provides robust feature representations that are more resilient to hard negatives. Additionally, the larger value helps mitigate the noise introduced by false negatives.
@@ 3.4 U-MARVEL Framework
Motivated by the findings presented above, we introduce a unified framework, term U-MARVEL (Universal MultimodAl RetrieVal via Embedding Learning), which consists of the following three stages: (1) Progressive transition: the model is progressively fine-tuned on the increasing complexity levels of the retrieval data, which enables the model to gradually adapt to retrieval tasks. (2) Hard Negative Mining and Fusion Reranker Model: building on Progressive transition, we first train a hard negative model for recall and a reranker model for ranking, and then linearly combine them to obtain a recall–reranker model. (3) Distillation: we perform improved knowledge distillation on the recall–reranker model. Detailed experimental settings and results for U-MARVEL are provided in Table 12 in Appendix C and Table 15 in Appendix D, respectively. Additionally, to demonstrate the generalization ability of our framework across different model sizes, we also provide results using the Qwen3-VL-4B-Instruct (Team, 2025) backbone in Table 16 and Appendix D.
@@ 4.1 M-BEIR Evaluation
We first compare U-MARVEL with competitors on M-BEIR benchmark (Wei et al., 2024) in both local and global pool settings, details of this dataset can be found in Appendix A. For fair comparison, we also present results of the recall-then-rerank cascade pipeline, denoted as , similar to that used in LamRA (Liu et al., 2024b), details seen in Appendix C. As shown in Table 7 and 17, U-MARVEL establishes a new state-of-the-art. Beyond the raw metrics, three key observations highlight the efficacy of our design: 1) Efficiency-performance trade-off: In the single-model setting, U-MARVEL significantly outperforms existing state-of-the-art approaches by a large margin. Furthermore, while comparisons with stronger baselines like LamRA (Liu et al., 2024b) typically involve a computationally expensive two-stage retrieval process, U-MARVEL achieves comparable performance using only single-stage inference. This dual advantage of superior accuracy and efficiency is directly attributable to our improved distillation strategy (Sec. 3.3), which successfully transfers the discriminative power of the reranker into the embedding space. 2) Robustness across settings: As detailed in Table 17 and Appendix D, our method maintains its superiority even in the more challenging “Global Pool” setting, where candidates from all tasks are mixed. It suggests that our method helps the model learn more robust, task-agnostic features rather than overfitting to specific task distributions. 3) Contribution of Components: The ablation study in Table 15 (Appendix C) further reveals that our model already surpasses existing state-of-the-art methods in the single-model setting upon the completion of “Progressive Transition”. Subsequently, “Hard Negative Mining” provides an additional boost in discrimination, while the “Distillation” stage is critical for closing the gap with reranking models.
| Methods | Avg. | ||||||||||||||||
| VN | CO | F200 | WQ | ES | WQ | VN | CO | F200 | NS | ON | InS | FQ | CR | ON | InS | ||
| R@5 | R@5 | R@10 | R@5 | R@5 | R@5 | R@5 | R@5 | R@10 | R@5 | R@5 | R@5 | R@10 | R@5 | R@5 | R@5 | ||
| Single model | |||||||||||||||||
| CLIP-L (Radford et al., 2021) | 43.3 | 61.1 | 6.6 | 36.2 | 43.3 | 45.1 | 41.3 | 79 | 7.7 | 26.1 | 24.2 | 20.5 | 7 | 13.2 | 38.8 | 26.4 | 32.5 |
| SigLIP (Zhai et al., 2023) | 30.1 | 75.7 | 36.5 | 39.8 | 27 | 43.5 | 30.8 | 88.2 | 34.2 | 28.9 | 29.7 | 25.1 | 14.4 | 22.7 | 41.7 | 27.4 | 37.2 |
| UniIR-BLIP (Wei et al., 2024) | 23.4 | 79.7 | 26.1 | 80 | 50.9 | 79.8 | 22.8 | 89.9 | 28.9 | 33 | 41 | 22.4 | 29.2 | 52.2 | 55.8 | 33 | 46.8 |
| UniIR-CLIP (Wei et al., 2024) | 42.6 | 81.1 | 18 | 84.7 | 59.4 | 78.7 | 43.1 | 92.3 | 18.3 | 32 | 45.5 | 27.9 | 24.4 | 44.6 | 67.6 | 48.9 | 50.6 |
| LamRA-Ret (Liu et al., 2024b) | 41.6 | 81.5 | 28.7 | 86 | 62.6 | 81.2 | 39.6 | 90.6 | 30.4 | 32.1 | 54.1 | 52.1 | 33.2 | 53.1 | 76.2 | 63.3 | 56.6 |
[…]
| +Reranker | |||||||||||||||||
| LamRA (Liu et al., 2024b) | 48 | 85.2 | 32.9 | 96.7 | 75.8 | 87.7 | 48.6 | 92.3 | 36.1 | 33.5 | 59.2 | 64.1 | 37.8 | 63.3 | 79.2 | 78.3 | 63.7 |
[…]
@@ 4.2 Zero-shot Evaluation
We evaluate the performance of U-MARVEL with several unseen datasets in a zero-shot manner. Details regarding the datasets and evaluation metrics are provided in Appendix B. As presented in Table 8, U-MARVEL achieves state-of-the-art performance on 9 out of 12 tasks, highlighting its exceptional zero-shot capabilities compared to leading methods such as VLM2Vec (Jiang et al., 2024b), UniME (Gu et al., 2025), and LamRA (Liu et al., 2024b). This verifies that our model effectively transfers knowledge from the diverse M-BEIR tasks to unseen domains, likely due to the high diversity of the progressive training curriculum that prevents catastrophic forgetting of the pre-trained knowledge. Furthermore, we evaluate U-MARVEL on the zero-shot text-to-video retrieval task. The results on the MSVD and MSR-VTT datasets, summarized in Table 9, demonstrate that U-MARVEL outperforms VLM2Vec (Jiang et al., 2024b), LamRA (Liu et al., 2024b), and LLaVE-7B (Lan et al., 2025). However, it is worth noting that exhibits degraded performance, which may be attributed to fail to model temporal contexts, such as action sequence progression.
| Methods | ITM | |||||||||||
| ST4V | Flickr | ST4V | Flickr | C-Neg | ||||||||
| R@1 | R@1 | R@1 | R@1 | R@1 | R@1 | MAP@5 | R@1 | R@1 | R@5 | Acc. | Acc. | |
| Single model | ||||||||||||
| CLIP-L (Radford et al., 2021) | 84 | 52.8 | 67.3 | 81.8 | 68.7 | 87.2 | 4 | 13.3 | 23.7 | 17.7 | 66.7 | 73 |
| (Wei et al., 2024) | 85.8 | 75 | 78.7 | 84.1 | 78.4 | 94.2 | 12.5 | 16.8 | 26.8 | 39.4 | 79.9 | 80.3 |
| E5-V (Jiang et al., 2024a) | 86.7 | 84 | 79.5 | 84 | 82.4 | 88.2 | 24.8 | 18.5 | 54.6 | 19.2 | 83.2 | 84.7 |
| MagicLens-L (Zhang et al., 2024b) | 85.5 | 59.3 | 72.5 | 60.9 | 24.2 | 84.6 | 29.6 | 16.3 | 28 | 22.6 | 62.7 | 75.9 |
| VLM2Vec (Jiang et al., 2024b) | 90.7 | 90.8 | 76 | 85.8 | 84.7 | 90.6 | - | - | - | - | - | 79.5 |
| UniME (Gu et al., 2025) | 97.2 | 95.9 | 81.9 | 93.9 | 95.2 | 93.4 | - | - | - | - | - | 85 |
| LamRA-Ret (Liu et al., 2024b) | 93.3 | 95.1 | 82.8 | 88.1 | 94.3 | 92.7 | 33.2 | 18.9 | 62.8 | 60.9 | 79.6 | 85.8 |
[…]
| +Reranker | ||||||||||||
| LamRA (Liu et al., 2024b) | 97.9 | 98.8 | 88.1 | 96.5 | 98 | 97.6 | 42.8 | 24.8 | 70.9 | 63.9 | 85.9 | 93.5 |
[…]
@@
| InternVideo (Wang et al., 2022) | 40.0 | 65.3 | 74.1 | 43.4 | 69.9 | 79.1 |
| ViCLIP (Wang et al., 2023) | 42.4 | - | - | 49.1 | - | - |
| UMT-L (Li et al., 2023b) | 42.6 | 64.4 | 73.1 | 49.9 | 77.7 | 85.3 |
| InternVideo2s2-6B (Wang et al., 2024b) | 55.9 | 78.3 | 85.1 | 59.3 | 84.4 | 89.6 |
@@
| VLM2Vec (Jiang et al., 2024b) | 43.5 | 69.3 | 78.9 | 49.5 | 77.5 | 85.7 |
| LamRA (Liu et al., 2024b) | 44.7 | 68.6 | 78.6 | 52.4 | 79.8 | 87.0 |
| LLaVE-7B (Lan et al., 2025) | 46.8 | 71.1 | 80.0 | 52.9 | 80.1 | 87.0 |
[…]